OpenStudy (anonymous):

Solve the equation 2/3 k + 6 = −12 for k. please provide work and answers

OpenStudy (stamp):

\[\frac{2}{3}k+6=-12\]This is a two-step algebra equation. Perform Step 1 and report back.

OpenStudy (anonymous):

i don't know how to thats why i posted it

OpenStudy (stamp):

Open your textbook girl and start reading, I 100% guarantee you there is an example and an explanation in there.

OpenStudy (anonymous):

i don't have a textbook i am home schooled

OpenStudy (stamp):

You do not have to be in public or private school to be in possession of a textbook. What is the source of your material?

OpenStudy (anonymous):

i do online school

OpenStudy (anonymous):

\[\frac{2}{3}k+6 = -12\]

OpenStudy (shaik0124):